OpenStudy (anonymous):

Write the equation of the quadratic function with roots -1 and 1 and a vertex at (0, -3).

OpenStudy (whpalmer4):

If you have a polynomial with a known set of roots \(r_1, r_2, ... r_n\) you can write the equation in factored form as \[P(x) = a(x-r_1)(x-r_2)...(x-r_n)\]where \(a\) is a constant (usually = 1) that allows you to make the curve pass through an arbitrary point (such as \((0,-3)\) in this case).

OpenStudy (anonymous):

Okay can you help me still I don't get that at all

OpenStudy (whpalmer4):

Well, what are your roots?

OpenStudy (anonymous):

I don't know what that means

OpenStudy (whpalmer4):

read the problem description again.

OpenStudy (anonymous):

-1 and 1

OpenStudy (whpalmer4):

"quadratic function with roots -1 and 1 "

OpenStudy (whpalmer4):

do you understand what a root is?

OpenStudy (anonymous):

Yes I do

OpenStudy (whpalmer4):

okay. it's a value of x that makes the polynomial have value 0, right?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

and the polynomial will be 0 at each root, and nowhere else, correct?

OpenStudy (anonymous):

I think so

OpenStudy (jdoe0001):

\(\bf x=-1\implies x+1=0\implies {\color{blue}{ (x+1)}}=0\\ \quad \\ x=1\implies x-1=0\implies {\color{blue}{ (x-1)}}=0\\ \quad \\ a{\color{blue}{ (x+1)}}{\color{blue}{ (x-1)}}=0\implies a{\color{blue}{ (x+1)}}{\color{blue}{ (x-1)}}=\textit{original polynomial}\)

OpenStudy (anonymous):

Omg make is easier

OpenStudy (whpalmer4):

that's correct. so, because a*b = 0 if either a or b = 0 (or both!), we can construct our polynomial from our roots by setting up product terms that will = 0 at each root. \[(x-1) = 0\]\[(x+1)=0\]those will equal 0 at x = 1, and x = -1

OpenStudy (whpalmer4):

Agreed?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Okay, so that means we can write our polynomial as\[P(x) = a(x+1)(x-1)\]All we have to do now is find the right value of \(a\) so that the polynomial goes through the point \((0,-3)\)

OpenStudy (whpalmer4):

Any idea how we might do that?

OpenStudy (anonymous):

No I don't

OpenStudy (whpalmer4):

Well, how about if I write it like this: \[y = P(x) = a(x+1)(x-1)\]\[(0,-3)\]

OpenStudy (whpalmer4):

we want to find \(a\) such that if we set \(x=0\), we get \(y = -3\)

OpenStudy (anonymous):

where do we plug in Y

OpenStudy (whpalmer4):

often these problems just ask for you to write a polynomial with the given roots. There's an infinite number of them, one for each possible value of \(a\). In that case, we just say \(a=1\) and are done. But in this case, we need 1 out of that infinity of choices, so we plug in our known point: \[y = P(x) = a(x+1)(x-1)\]\[-3 = P(0) = a(0+1)(0-1)\]\[-3 = a(1)(-1)\]

OpenStudy (whpalmer4):

What do you get for \(a\)?

OpenStudy (anonymous):

a= 3?

OpenStudy (whpalmer4):

very good! So what is our polynomial?

OpenStudy (anonymous):

f(x) = -3x^2 − 3x

OpenStudy (anonymous):

?

OpenStudy (whpalmer4):

mmm...no, not quite. try again?

OpenStudy (anonymous):

f(x) = 3x^2 + 3

OpenStudy (whpalmer4):

getting closer :-)

OpenStudy (anonymous):

f(x) = 3x^2 − 3

OpenStudy (whpalmer4):

Bingo!

OpenStudy (anonymous):

can you keep helping me ? I have 9 more question

OpenStudy (whpalmer4):

Here's a plot of that for printing and hanging on the wall after you verify that it crosses the x-axis at x = 1, x = -1, and passes through (0,-3):

OpenStudy (anonymous):

can you keep helping me ? I have 9 more question

OpenStudy (whpalmer4):

Yes, but before we go on, I want to beat this horse to death a little bit more :-) Remember I said there are an infinite number of polynomials that have that same set of roots? Here I've plotted a number of different values of \(a\) (1,2,3,-2). Notice how they all have the same roots, even though the rest of the curve varies?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you going to help me or not

OpenStudy (whpalmer4):

patience, it's not like I'm getting paid to do this! I'll be back in a few minutes.

OpenStudy (anonymous):

promise c;

OpenStudy (whpalmer4):

Okay, the others are finally out of the house, you have my full attention :-)

OpenStudy (anonymous):

Find the equation of the quadratic function with zeros 10 and 12 and a vertex at (11, -2). HELP

OpenStudy (whpalmer4):

Okay, this is exactly what we just did. zeros (aka roots) are x = 10, x = 12. What is our equation?

OpenStudy (anonymous):

uh I don't know

OpenStudy (anonymous):

I think the answer is y = 2x^2 − 44x − 240

OpenStudy (whpalmer4):

Okay, here's the drill again. To construct a polynomial \(P(x)\) with a given set of roots (or zeros), \(r_1,r_2,...r_n\) \[P(x) = a(x-r_1)(x-r_2)...(x-r_n)\]If the polynomial must pass through a given point \((x_0,y_0)\), set \[P(x_0) = y_0\]and solve for the necessary value of \(a\), otherwise let \(a = 1\).

OpenStudy (anonymous):

I think the answer y = 2x^2 − 44x − 240

OpenStudy (whpalmer4):

Uh, you have an incorrect sign in there, it appears.

OpenStudy (anonymous):

y = 2x^2 − 44x + 240

OpenStudy (whpalmer4):

Two tests to see if you've gotten a correct answer (both must be satisfied): 1) evaluating the polynomial at the zeros must give you 0 as a result 2) the polynomial must pass through the target point \[y = 2x^2-44x+240\]\[2(10)^2-44(10)+240 =200-440+240 = 0\checkmark\]\[2(12)^2-44(12)+240 = 288-528+240 = 0\checkmark\]\[-2 = 2(11)^2-44(11)+240\]\[-2=242-484+240\checkmark\]